# A Setwise Approach for Effective and Highly Efficient Zero-shot Ranking with Large Language Models

### v1

@@ Abstract.
Large Language Models (LLMs) demonstrate impressive effectiveness in zero-shot document ranking tasks. Pointwise, Pairwise, and Listwise prompting approaches have been proposed for LLM-based zero-shot ranking. Our study begins by thoroughly evaluating these existing approaches within a consistent experimental framework, considering factors like model size, token consumption, latency, among others. This first-of-its-kind comparative evaluation of these approaches allows us to identify the trade-offs between effectiveness and efficiency inherent in each approach. We find that while Pointwise approaches score high on efficiency, they suffer from poor effectiveness. Conversely, Pairwise approaches demonstrate superior effectiveness but incur high computational overhead. To further enhance the efficiency of LLM-based zero-shot ranking, we propose a novel Setwise prompting approach. Our approach reduces the number of LLM inferences and the amount of prompt token consumption during the ranking procedure, significantly improving the efficiency of LLM-based zero-shot ranking. We test our method using the TREC DL datasets and the BEIR zero-shot document ranking benchmark. The empirical results indicate that our approach considerably reduces computational costs while also retaining high zero-shot ranking effectiveness.
@@ 1. Introduction
Large Language Models (LLMs) such as GPT-3 (Brown et al., 2020), FlanT5 (Wei et al., 2021), and PaLM (Chowdhery et al., 2022) have been shown highly effective across a diverse range of natural language processing tasks under the zero-shot settings (Kojima et al., 2022; Brown et al., 2020; Agrawal et al., 2022; Wang et al., 2023). Notably, these LLMs have also been adapted for zero-shot document ranking tasks, exhibiting strong zero-shot ranking capabilities (Liang et al., 2022; Sachan et al., 2022; Sun et al., 2023; Ma et al., 2023; Pradeep et al., 2023; Qin et al., 2023). The methodologies for harnessing LLMs in zero-shot ranking tasks can be broadly categorized into three main approaches: Pointwise (Liang et al., 2022; Sachan et al., 2022), Listwise (Sun et al., 2023; Ma et al., 2023; Pradeep et al., 2023), and Pairwise (Qin et al., 2023). These approaches employ different prompting strategies to instruct the LLM to output a relevance estimation for each candidate document. While these LLM-based zero-shot ranking approaches have been successful individually, it is worth noting that there has been a lack of fair comparison in the literature regarding their effectiveness, and in particular, their efficiency within the exact same experimental framework. This includes factors such as utilizing the same size of LLM, evaluation benchmarks, and computational resources. We believe it is very important to establish a rigorous framework for evaluating these LLM-based zero-shot ranking approaches. By doing so, we can draw meaningful conclusions about their comparative effectiveness and efficiency.
@@
Thus, in this paper, we first conduct a systematic evaluation of all existing approaches within a consistent experimental environment. In addition to assessing ranking effectiveness, we also compare the efficiency of these methods in terms of computational expenses and query latency. Our findings indicate that the Pairwise approach emerges as the most effective but falls short in terms of efficiency even with the assistance of sorting algorithms aimed at improving this. Conversely, the Pointwise approach stands out as the most efficient but lags behind other methods in terms of ranking effectiveness. The Listwise approach, which relies solely on the generation of document labels in order, can strike a middle ground between efficiency and effectiveness but this varies considerably based on configuration, implementation and evaluation dataset (highlighting the importance of thoroughly evaluating these model under multiple settings). Overall, these comprehensive results furnish an understanding of the strengths and weaknesses of these LLM-based zero-shot ranking approaches, providing valuable insights for practitioners seeking to select the most suitable approach for real-world applications.
@@
Having considered all the different approaches and their results in terms of efficiency and effectiveness tradeoffs, we set about devising a method that was both effective and efficient. Our approach was to take the most effective model (Pairwise) and to enhance its efficiency (without seriously compromising effectiveness). Our solution is a novel Setwise prompting approach. This concept stems from our realization that the sorting algorithms employed by Pairwise approaches can be accelerated by comparing multiple documents, as opposed to just a pair at a time.
@@
Our Setwise prompting approach instructs LLMs to select the most relevant document to the query from a set of candidate documents. This straightforward adjustment allows the sorting algorithms to infer relevance preferences for more than two candidate documents at each step, thus significantly reducing the total number of comparisons required; this leads to substantial savings in computational resources. Furthermore, beyond the adjustment to Pairwise approaches, Setwise prompting allows the utilization of model output logits to estimate the likelihood of ranks of document labels, a capability not feasible in existing Listwise approaches, which solely rely on document label ranking generation —– a process that is slow and less effective. We evaluate our Setwise approach along with other existing approaches under the same experimental setting. Our results show that the incorporation of our Setwise prompting substantially improves the efficiency of both Pairwise and Listwise approaches. In addition, Setwise sorting enhances Pairwise and Listwise robustness to variations in the internal ordering quality of the initial rankings: no matter what the initial ordering of the top-k documents to rank is, our method provides consistent and effective results. This is unlike other methods that are highly susceptible to such initial ordering.
@@ 3.1. Limitations of Current Approaches
3.2. Speeding-up Pairwise with Setwise
With our prompt, sorting-based Pairwise approaches can be considerably accelerated. This is because the original heap sort and bubble sort algorithm used in the Pairwise approach only compares a pair of documents at each step in the sorting process, as illustrated in Figure 1(a) and 1(c). These sorting algorithms can be sped up by comparing more than two documents at each step. For example, in the heap sort algorithm, the “heapify" function needs to be invoked for each subtree, where the parent node must be swapped with the child node with the highest value if it exceeds the parent value. In the case of Figure 1(a), to perform “heapify" with pairwise prompting, a minimum of 6 comparisons (each root node paired with each child node) are required. Conversely, if we increase the number of child nodes in each subtree to 3 and can compare 4 nodes at a time, only 2 comparisons are needed to “heapify" a tree with 9 nodes, as illustrated in Figure 1(b). Similarly, for the bubble sort algorithm, if we can compare more than a pair of documents at a time, each “bubbling” process will be accelerated. For instance, in Figure 1(c), there are 4 comparisons in total, but in Figure 1(d), with the ability to compare 3 documents at once, only 2 comparisons are required to be able to bring the node with the largest value to the top. Our Setwise prompting is designed to instruct LLMs to compare the relevance of multiple documents at a time, making it well-suited for this purpose.
3.3. Listwise Likelihoods with Setwise
@@ 4.1. Datasets and evaluations
Flan-t5-large | pointwise.qlm | .557 | 100 | 15211.6 | - | 0.6 | .567 | 100 | 15285.2 | - | 0.5 | |
pointwise.yes_no | .654 | 100 | 16111.6 | - | 0.6 | .615 | 100 | 16185.2 | - | 0.6 | ||
listwise.generation | .561 | 245 | 119120.8 | 2581.35 | 54.2 | .547 | 245 | 119629.6 | 2460.1 | 52.0 | ||
listwise.likelihood | .669 | 245 | 94200.7 | - | 10.0 | .626 | 245 | 95208.3 | - | 10.0 | ||
pairwise.allpair | .666 | 9900 | 3014383.1 | 49500 | 109.6 | .622 | 9900 | 3014232.7 | 49500 | 108.9 | ||
pairwise.heapsort | .657 | 230.3 | 104952.5 | 2303.3 | 16.1 | .619 | 226.8 | 104242.1 | 2268.3 | 16.1 | ||
pairwise.bubblesort | .636 | 844.2 | 381386.3 | 8441.6 | 58.3 | .589 | 778.5 | 357358.5 | 7785.4 | 54.1 | ||
setwise.heapsort | .670 | 125.4 | 40460.6 | 626.9 | 8.0 | .618 | 124.2 | 40362.0 | 621.1 | 8.0 | ||
setwise.bubblesort | .678 | 460.5 | 147774.1 | 2302.3 | 29.1 | .624 | 457.4 | 148947.3 | 2287.1 | 28.9 | ||
Flan-t5-xl | pointwise.qlm | .542 | 100 | 15211.6 | - | 1.4 | .542 | 100 | 15285.2 | - | 1.4 | |
pointwise.yes_no | .650 | 100 | 16111.6 | - | 1.5 | .636 | 100 | 16185.2 | - | 1.5 | ||
listwise.generation | .569 | 245 | 119163.0 | 2910 | 71.4 | .547 | 245 | 119814.3 | 2814.7 | 69.0 | ||
listwise.likelihood | .689 | 245 | 94446.1 | - | 12.5 | .672 | 245 | 95298.7 | - | 12.6 | ||
pairwise.allpair | .713 | 9900 | 2953436.2 | 49500 | 254.9 | .682 | 9900 | 2949457.6 | 49500 | 254.8 | ||
pairwise.heapsort | .705 | 241.9 | 110126.9 | 2418.6 | 20.5 | .692 | 244.3 | 111341 | 2443.3 | 20.8 | ||
pairwise.bubblesort | .683 | 886.9 | 400367.1 | 8869.1 | 75.1 | .662 | 863.9 | 394954.2 | 8638.5 | 74.3 | ||
setwise.heapsort | .693 | 129.5 | 41665.7 | 647.4 | 9.6 | .678 | 127.8 | 41569.1 | 638.9 | 9.7 | ||
setwise.bubblesort | .705 | 466.9 | 149949.1 | 2334.5 | 35.2 | .676 | 463.5 | 151249.8 | 2317.6 | 35.3 | ||
Flan-t5-xxl | pointwise.qlm | .506 | 100 | 15211.6 | - | 3.7 | .492 | 100 | 15285.2 | - | 3.7 | |
pointwise.yes_no | .644 | 100 | 16111.6 | - | 3.9 | .632 | 100 | 16185.2 | - | 3.9 | ||
listwise.generation | .662 | 245 | 119334.7 | 2824 | 100.1 | .637 | 245 | 119951.6 | 2707.9 | 97.3 | ||
listwise.likelihood | .701 | 245 | 94537.5 | - | 36.6 | .690 | 245 | 95482.7 | - | 36.9 | ||
pairwise.allpair | .699 | 9900 | 2794942.6 | 49500 | 730.2 | .688 | 9900 | 2794928.4 | 49500 | 730.5 | ||
pairwise.heapsort | .708 | 239.4 | 109402 | 2394 | 45.0 | .699 | 240.5 | 110211.8 | 2404.8 | 45.2 | ||
h | pairwise.bubblesort | .679 | 870.5 | 394386 | 8705.3 | 162.5 | .681 | 842.9 | 387359.2 | 8428.5 | 158.8 | |
i | setwise.heapsort | .706 | 130.1 | 42078.6 | 650.5 | 20.2 | .688 | 128.1 | 41633.7 | 640.6 | 20.0 | |
setwise.bubblesort | .711 | 468.3 | 150764.8 | 2341.6 | 72.6 | .686 | 467.9 | 152709.5 | 2339.6 | 73.2 |
@@ 4.2. Implementation details
To establish the initial BM25 first-stage ranking for all datasets, we employed the Pyserini Python library (Lin et al., 2021) with default settings. For LLM-based zero-shot re-rankers, we followed the prompts recommended in existing literature to guide Flan-t5 models of varying sizes (Flan-t5-large with 780M parameters, Flan-t5-xl with 3B parameters, and Flan-t5-xxl with 11B parameters) in executing the zero-shot ranking task.

### v2

@@ Abstract.
We propose a novel zero-shot document ranking approach based on Large Language Models (LLMs): the Setwise prompting approach. Our approach complements existing prompting approaches for LLM-based zero-shot ranking: Pointwise, Pairwise, and Listwise. Through the first-of-its-kind comparative evaluation within a consistent experimental framework and considering factors like model size, token consumption, latency, among others, we show that existing approaches are inherently characterised by trade-offs between effectiveness and efficiency. We find that while Pointwise approaches score high on efficiency, they suffer from poor effectiveness. Conversely, Pairwise approaches demonstrate superior effectiveness but incur high computational overhead. Our Setwise approach, instead, reduces the number of LLM inferences and the amount of prompt token consumption during the ranking procedure, compared to previous methods. This significantly improves the efficiency of LLM-based zero-shot ranking, while also retaining high zero-shot ranking effectiveness. We make our code and results publicly available at https://github.com/ielab/llm-rankers.
@@ 1. Introduction
Large Language Models (LLMs) such as GPT-3 (Brown et al., 2020), FlanT5 (Wei et al., 2021), Gemini (Team et al., 2023), LLaMa2 (Touvron et al., 2023b), and PaLM (Chowdhery et al., 2022) are highly effective across a diverse range of natural language processing tasks under the zero-shot settings (Kojima et al., 2022; Brown et al., 2020; Agrawal et al., 2022; Wang et al., 2023). Notably, these LLMs have also been adapted for zero-shot document ranking. (Liang et al., 2022; Sachan et al., 2022; Sun et al., 2023; Ma et al., 2023; Pradeep et al., 2023; Qin et al., 2023). Using LLMs in zero-shot ranking tasks can be broadly categorized into three main approaches: Pointwise (Liang et al., 2022; Sachan et al., 2022; Zhuang et al., 2023), Listwise (Sun et al., 2023; Ma et al., 2023; Pradeep et al., 2023), and Pairwise (Qin et al., 2023). These approaches employ different prompting strategies to instruct the LLM to output a relevance estimation for each candidate document and rank documents accordingly. Compared to traditional neural ranking methods (Lin et al., 2021b), these LLM-based rankers do not require any further supervised fine-tuning and exhibit strong zero-shot ranking capabilities. While these LLM-based zero-shot ranking approaches have been successful individually, it is worth noting that there has been a lack of fair comparison in the literature regarding their effectiveness, and in particular, their efficiency within the exact same experimental framework. This includes factors such as utilizing the same size LLM, evaluation benchmarks, and computational resources. We believe it is critical to establish a rigorous framework for evaluating these LLM-based zero-shot ranking approaches. By doing so, we can draw meaningful conclusions about their comparative effectiveness and efficiency.
@@
Thus, in this paper, we first conduct a systematic evaluation of all existing approaches within a consistent experimental environment. In addition to assessing ranking effectiveness, we also compare the efficiency of these methods in terms of computational expense and query latency. Our findings indicate that the Pairwise approach emerges as the most effective but falls short in terms of efficiency even with the assistance of sorting algorithms aimed at improving efficiency. Conversely, the Pointwise approach stands out as the most efficient but lags behind other methods in terms of ranking effectiveness. The Listwise approach, which relies solely on the generation of document labels in order, can strike a middle ground between efficiency and effectiveness but this varies considerably based on configuration, implementation and evaluation dataset (highlighting the importance of thoroughly evaluating these model under multiple settings). Overall, these comprehensive results offer an understanding of the strengths and weaknesses of LLM-based zero-shot ranking approaches, providing valuable insights for those seeking to select the most suitable approach for real-world applications.
@@
Having considered all the different approaches and their results in terms of efficiency and effectiveness tradeoffs, we set about devising a method that was both effective and efficient. Our approach was to take the most effective model (Pairwise) and to enhance its efficiency (without seriously compromising effectiveness). Our solution is a novel Setwise prompting approach. This concept stems from our realisation that the sorting algorithms employed by Pairwise approaches can be accelerated by comparing multiple documents, as opposed to just a pair at a time.
@@
Our Setwise prompting approach instructs LLMs to select the most relevant document to the query from a set of candidate documents. This straightforward adjustment allows the sorting algorithms to infer relevance preferences for more than two candidate documents at each step, thus significantly reducing the total number of comparisons required; this leads to substantial savings in computational resources. Furthermore, beyond the adjustment to Pairwise approaches, Setwise prompting allows for the utilization of model output logits to estimate the likelihood of ranks of document labels, a capability not feasible in existing Listwise approaches, which solely rely on document label ranking generation —– a process that is slow and less effective. We Setwise and other existing approaches under the same experimental settings to provide a clear and consistent comparison. Our results show that the incorporation of our Setwise prompting substantially improves the efficiency of both Pairwise and Listwise approaches. In addition, Setwise sorting enhances Pairwise and Listwise robustness to variations in the internal ordering quality of the initial rankings: no matter what the initial ordering of the top-k documents to rank is, our method provides consistent and effective results. This is unlike other methods that are highly susceptible to such initial ordering.
@@ 2.3. Pairwise prompting approaches
2.4. Other Directions in using LLMs for Ranking
The three families of approaches outlined above directly use prompting of LLMs to infer document relevance to a given query, leveraging LLMs for document re-ranking in a zero-shot setting with no training data required for the ranking task. An alternative direction in using LLMs for retrieval has also emerged, where LLMs are instead used as text embedding models for dense document retrieval (Lee et al., 2024; Wang et al., 2024a, b; BehnamGhader et al., 2024). Because these methods create document representations independently of query representations, they can be used as a first-stage document retriever (i.e. in all effects as a bi-encoder architecture), rather than being limited to be used as a re-ranker as instead are the approaches from the three families we have reviewed above. However, these methods require performing contrastive learning training to enable the generative LLM to act as a text-embedding model. An exception to this is the recent PromptReps method (Zhuang et al., 2024), which does not require contrastive training, achieving document (and query) embedding simply via prompt engineering. Another feature of PromptReps is the ability to obtain at the same time both a dense and a sparse representation of documents and queries.
@@ 3. Setwise Ranking Prompting
In this section, we discuss the limitations present in the current LLM-based zero-shot ranking methods. We then describe our proposed Setwise approach and how it addresses these limitations.
@@ 3.1. Limitations of Current Approaches
Next, we introduce our new Setwise prompting approaches, designed to overcome the efficiency limitations of both Pairwise and Listwise methods by minimizing the number of required LLM inferences and leveraging the logits produced by the LLM.
@@
3.2. Speeding-up Pairwise with Setwise
With our prompt, sorting-based Pairwise approaches can be considerably accelerated. This is because the original heap sort and bubble sort algorithm used in the Pairwise approach only compares a pair of documents at each step in the sorting process, as illustrated in Figure 2(a) and 2(c). These sorting algorithms can be sped up by comparing more than two documents at each step. For example, in the heap sort algorithm, the “heapify" function needs to be invoked for each subtree, where the parent node must be swapped with the child node with the highest value if it exceeds the parent value. In the case of Figure 2(a), to perform “heapify" with pairwise prompting, a minimum of 6 comparisons (each root node paired with each child node) are required. Conversely, if we increase the number of child nodes in each subtree to 3 and can compare 4 nodes at a time, only 2 comparisons are needed to “heapify" a tree with 9 nodes, as illustrated in Figure 2(b). Similarly, for the bubble sort algorithm, if we can compare more than a pair of documents at a time, each “bubbling” process will be accelerated. For instance, in Figure 2(c), there are 4 comparisons in total, but in Figure 2(d), with the ability to compare 3 documents at once, only 2 comparisons are required to be able to bring the node with the largest value to the top. Our Setwise prompting is designed to instruct LLMs to compare the relevance of multiple documents at a time, making it well-suited for this purpose.
3.3. Listwise Likelihoods with Setwise
@@ 3.4. Advantages of Setwise
On the other hand, model output calibration might be a concern for Pointwise methods because each document’s relevance is inferred independently. Consequently, ranking documents based on Pointwise relevance scores necessitates calibration (Qin et al., 2023). However, for our Setwise method (as well as for Pairwise and Listwise) the model output logits are derived from the input of multiple documents, which do not directly represent the relevance of a single document but rather serve as an indicator of preference among documents. Thus, calibration is not necessary.
@@ 4.1. Datasets and evaluations
Flan-t5-large |
pointwise.qlm | .557 | 100 | 15211.6 | - | 0.6 | .567a | 100 | 15285.2 | - | 0.5 | |
pointwise.yes_no | .654abd | 100 | 16111.6 | - | 0.6 | .615ad | 100 | 16185.2 | - | 0.6 | ||
listwise.generation | .561a | 245 | 119120.8 | 2581.35 | 54.2 | .547a | 245 | 119629.6 | 2460.1 | 52 | ||
listwise.likelihood | .669abd | 245 | 94200.7 | - | 10 | .626abd | 245 | 95208.3 | - | 10 | ||
pairwise.allpair | .666abd | 9900 | 3014383.1 | 49500 | 109.6 | .622abd | 9900 | 3014232.7 | 49500 | 108.9 | ||
pairwise.heapsort | .657abd | 230.3 | 104952.5 | 2303.3 | 16.1 | .619abd | 226.8 | 104242.1 | 2268.3 | 16.1 | ||
pairwise.bubblesort | .636abd | 844.2 | 381386.3 | 8441.6 | 58.3 | .589ad | 778.5 | 357358.5 | 7785.4 | 54.1 | ||
setwise.heapsort | .670abd | 125.4 | 40460.6 | 626.9 | 8.0 | .618ad | 124.2 | 40362.0 | 621.1 | 8.0 | ||
setwise.bubblesort | .678abdh | 460.5 | 147774.1 | 2302.3 | 29.1 | .624abd | 457.4 | 148947.3 | 2287.1 | 28.9 | ||
Flan-t5-xl |
pointwise.qlm | .542 | 100 | 15211.6 | - | 1.4 | .542a | 100 | 15285.2 | - | 1.4 | |
pointwise.yes_no | .650abd | 100 | 16111.6 | - | 1.5 | .636abd | 100 | 16185.2 | - | 1.5 | ||
listwise.generation | .569a | 245 | 119163.0 | 2910 | 71.4 | .547a | 245 | 119814.3 | 2814.7 | 69 | ||
listwise.likelihood | .689abd | 245 | 94446.1 | - | 12.5 | .672abd | 245 | 95298.7 | - | 12.6 | ||
pairwise.allpair | .713abcdehi | 9900 | 2953436.2 | 49500 | 254.9 | .682abcd | 9900 | 2949457.6 | 49500 | 254.8 | ||
pairwise.heapsort | .705abcd | 241.9 | 110126.9 | 2418.6 | 20.5 | .692abcdh | 244.3 | 111341 | 2443.3 | 20.8 | ||
pairwise.bubblesort | .683abd | 886.9 | 400367.1 | 8869.1 | 75.1 | .662abd | 863.9 | 394954.2 | 8638.5 | 74.3 | ||
setwise.heapsort | .693abcd | 129.5 | 41665.7 | 647.4 | 9.6 | .678abcd | 127.8 | 41569.1 | 638.9 | 9.7 | ||
setwise.bubblesort | .705abcd | 466.9 | 149949.1 | 2334.5 | 35.2 | .676abcd | 463.5 | 151249.8 | 2317.6 | 35.3 | ||
Flan-t5-xxl |
pointwise.qlm | .506 | 100 | 15211.6 | - | 3.7 | .492 | 100 | 15285.2 | - | 3.7 | |
pointwise.yes_no | .644ab | 100 | 16111.6 | - | 3.9 | .632ab | 100 | 16185.2 | - | 3.9 | ||
listwise.generation | .662ab | 245 | 119334.7 | 2824 | 100.1 | .637ab | 245 | 119951.6 | 2707.9 | 97.3 | ||
listwise.likelihood | .701abcd | 245 | 94537.5 | - | 36.6 | .690abcd | 245 | 95482.7 | - | 36.9 | ||
pairwise.allpair | .699abcd | 9900 | 2794942.6 | 49500 | 730.2 | .688abcd | 9900 | 2794928.4 | 49500 | 730.5 | ||
pairwise.heapsort | .708abcdh | 239.4 | 109402 | 2394 | 45 | .699abcd | 240.5 | 110211.8 | 2404.8 | 45.2 | ||
h | pairwise.bubblesort | .679ab | 870.5 | 394386 | 8705.3 | 162.5 | .681abcd | 842.9 | 387359.2 | 8428.5 | 158.8 | |
i | setwise.heapsort | .706abcd | 130.1 | 42078.6 | 650.5 | 20.2 | .688abcd | 128.1 | 41633.7 | 640.6 | 20.0 | |
setwise.bubblesort | .711abcdh | 468.3 | 150764.8 | 2341.6 | 72.6 | .686abcd | 467.9 | 152709.5 | 2339.6 | 73.2 |
@@ 4.2. Implementation details
To establish the initial BM25 first-stage ranking for all datasets, we employed the Pyserini Python library (Lin et al., 2021a) with default settings. For LLM-based zero-shot re-rankers, we followed the prompts recommended in existing literature to guide Flan-t5 models of varying sizes (Flan-t5-large with 780M parameters, Flan-t5-xl with 3B parameters, and Flan-t5-xxl with 11B parameters) in executing the zero-shot ranking task.
@@
BM25 | .595 | .322 | .442 | .318 | .679 | .331 | .395 | .407 | .436 | ||
Flan-t5-large |
pointwise.qlm | .664a | .322 | .260 | .305 | .644c | .314c | .413c | .439af | .420 | |
pointwise.yes_no | .664a | .308 | .238 | .296 | .504 | .275 | .346 | .456af | .386 | ||
listwise.generation | .692a | .333ac | .441bcefhi | .391abc | .650c | .343ace | .428ac | .441af | .465 | ||
listwise.likelihood | .756abcd | .334c | .327bc | .444abcdfgh | .639c | .308c | .453ac | .475abdfg | .467 | ||
pairwise.heapsort | .761abcdg | .336c | .318bc | .414abcd | .671chi | .325c | .440ac | .402 | .458 | ||
pairwise.bubblesort | .714a | .341abcdh | .447bcefhi | .416abcd | .700bcdefhi | .361abcdefh | .440ac | .439af | .482 | ||
setwise.heapsort | .768abcdg | .325c | .303c | .413abcd | .620c | .319c | .439c | .462abf | .456 | ||
setwise.bubblesort | .761abcdg | .338abch | .394bcefh | .441abcdfgh | .636c | .351bcefh | .447ac | .497abcdefgh | .483 | ||
Flan-t5-xl |
pointwise.qlm | .679a | .330 | .216 | .310c | .696c | .299 | .422 | .427 | .422 | |
pointwise.yes_no | .698a | .331 | .269 | .273 | .553 | .297 | .413 | .479ab | .414 | ||
listwise.generation | .650a | .334a | .451bcefghi | .366abc | .694c | .349abcefh | .437a | .475ab | .470 | ||
listwise.likelihood | .736abd | .360abcd | .310b | .449abcdfghi | .686c | .320 | .472abc | .526abcd | .482 | ||
pairwise.heapsort | .778abcde | .355abcd | .303b | .417abcd | .711ch | .317 | .471abc | .550abcdehi | .488 | ||
pairwise.bubblesort | .763abcd | .359abcd | .400bcefhi | .432abcdf | .734abcdefhi | .353bcefh | .485abcd | .553abcdehi | .510 | ||
setwise.heapsort | .757abcd | .352abcd | .283b | .428abcdf | .677c | .314 | .465ac | .520abcd | .475 | ||
setwise.bubblesort | .756abcd | .353abcd | .330bch | .438abcdfh | .691c | .362abcefh | .497abcd | .537abcdh | .496 | ||
Flan-t5-xxl |
pointwise.qlm | .707a | .342ac | .188 | .324 | .712c | .307c | .431 | .440a | .431 | |
pointwise.yes_no | .691a | .322 | .240b | .305 | .623 | .274 | .392 | .515ab | .420 | ||
listwise.generation | .664a | .344ac | .453bcefhi | .441abcefghi | .736ac | .353bcefh | .458ac | .495ab | .493 | ||
listwise.likelihood | .749acd | .352ac | .307bc | .416abch | .725ac | .316c | .479abc | .518abd | .483 | ||
pairwise.heapsort | .738acd | .359abcdhi | .324bc | .407abc | .744abc | .328c | .487abc | .543abcdeh | .491 | ||
pairwise.bubblesort | .733ad | .363abcdehi | .423bcefh | .421abcfh | .756abcdeh | .355bcefh | .490abcd | .550abcdehi | .511 | ||
setwise.heapsort | .752abcd | .346ac | .297bc | .402abc | .726ac | .321c | .473abc | .513ab | .479 | ||
setwise.bubblesort | .768abcdfg | .346ac | .388bcefh | .424abcfh | .754abceh | .343bceh | .479abc | .534abdeh | .505 |